OpenStudy (kkutie7):

match solutions and differentials equations. Note: each equation may have more than one solution. Select all that apply. I have three of them. The first: (a)\[3y"-3y=0\] \[y=e^{x}\],\[y=x^{3}\],\[y=e^{-x}\], or \[y=x^{-2}\] I think its easier for me to write a our like this: \[3\frac{d^{2}y}{dx^{2}}-3y=0\]

OpenStudy (kkutie7):

Also I want help not answers =)

OpenStudy (lochana):

I was reading about this a moment ago:). In homogeneous differential equations, we have this way of thinking as a quadratic equation. so what we need here is to make the differential equation into quadratic equation

OpenStudy (lochana):

assume \( \frac{d^2y}{dx^2} = m^2 \\ \frac{dy}{dx} = m^1 \\ y = m^0 = 1 \)

OpenStudy (kkutie7):

so I have to do integrals then to get it roughly to \[y=ax^{2}+bx+c\]

OpenStudy (lochana):

mm.. not really. we need to take roots from converted equation. And take solutions. Solutions are going to be different with the nature of roots (complex, non-complex etc)

OpenStudy (lochana):

I can solve this with steps. Do you want to see it?

OpenStudy (kkutie7):

yes I think I need to see it because nothing is clicking

OpenStudy (superdavesuper):

dont mean to butt in....but since the Q has provided choices for ans, wouldnt it be easier to simply differentiate twice each choice n substitute back into the DE to see if it works?

OpenStudy (lochana):

yes. that's right. I was trying to solve it in general way.:). you can do that too

OpenStudy (superdavesuper):

@lochana u can go ahead n show the general way :) im sure @Kkutie7 can do the differentiation on her own if she wanna try it!

OpenStudy (lochana):

so. do substitution, you get \(3m^2 - 3 = 0 \\ m^2 -1 \\ (m-1)(m+1)\ = 0 \\ m1= 1\ and\ m2 = -1 \), as you can see here, roots are real numbers. that means the solution should be in the form of \(y = Ae^{m1x} + Be^{m2x} \\ y = Ae^{-x}+Be^{x} \)

OpenStudy (lochana):

in you question, I see 2 solutions. \(y = e^{x} \ and \ y =e^{-x}\)

OpenStudy (lochana):

and check back substitution like @superdavesuper said

OpenStudy (lochana):

@Kkutie7 did you get it?

OpenStudy (lochana):

you can change "A" and "B" into any value. in your case, it is A = 1 ; B =1

OpenStudy (kkutie7):

yes i got it thank you

OpenStudy (lochana):

you are welcome:)